f(n)= 3*f(n/2) - 4 when n is even and f(1) = -3.
a. What is the value of f(4)?

b. What is the value of f(8)?


Answer by jim_thompson5910(35256)

I'll do part a) and leave you part b)

a)

f(n)= 3*f(n/2) - 4

f(2)= 3*f(2/2) - 4

f(2)= 3*f(1) - 4

f(2)= 3*(-3) - 4

f(2)= -13

---------------

f(n)= 3*f(n/2) - 4

f(4)= 3*f(4/2) - 4

f(4)= 3*f(2) - 4

f(4)= 3*(-13) - 4

f(4) = -43